# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v1

@@ Abstract
In this work, we study the gravitating lensing Euler-Heisenberg black hole surrounded by perfect fluid dark matter. This kind of black hole solution enables us to investigate the nontrivial interplay between the dark matter effects and nonlinear electrodynamics effects (or quantum electrodynamics effects). The important observable in gravitational lensing observations are calculated and discussed, including the gravitational deflection angle of light, time delay of light, precession angle of massive orbit’s bound orbits, and black hole shadow radius. The results indicates that the Euler-Heisenberg black hole with a negative perfect fluid dark matter parameter ($\lambda_{\text{DM}}<0$) could greatly intensify the gravitational deflection angle of light, time delay of light and precession angle of massive object’s bound orbit, while the nonlinear electrodynamics effects do not have large influences on these gravitational lensing observables.
Keywords: Euler Heisenberg Black Hole; Gravitational Lensing; Perfect Fluid Dark Matter;
Nonlinear Electrodynamics
@@ 1 Introduction
The equation of state for dark matter assembled in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted large interests [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trial effects and characters for dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and the effective spacetime metrics for several kinds of black holes (such as Schwarzschild, RN, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties are extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
Inspired by the simplicity and success of PFDM model in black hole physics and astronomy, we choose a novel black hole solution surrounded by PFDM — the Euler-Heisenberg black hole. The Euler-Heisenberg black hole, which is obtained from the Euler-Heisenberg effective action [53, 54, 55, 56], stimulated numerous studies in past few years, for its nontrivial realization on the coupling between nonlinear electrodynamics and gravity systems, especially with the quantum loop-corrections in quantum electrodynamics (QED) [57]. With these desired properties, this kind of black hole is often used to explore the non-linear electrodynamics in black hole physics, resulting in a number of studies on thermodynamic transitions, thermodynamic topology, black hole shadow, gravitational lensing, quasi-normal modes for Euler-Heisenberg black hole in recent years [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 82, 78, 79, 80, 81, 86, 83, 84, 88, 85, 87]. The nonlinear electrodynamics effects (or QED effects) on gravitation are extremely important issues in high energy scale, and they may have potential influences on the evolution of our universe (especially at the beginning stage), such as the inflationary theory in Big Bang, the dynamic evolution of dark matter or dark energy, and the cosmological accelerating expansion of universe. Therefore, it is interesting to study the interplay between dark matter and non-linear electrodynamics in gravitational systems. The Euler-Heisenberg black hole surrounded by PFDM could provide us a simply way to investigate their influences on charged black holes. Recently, several analytical solution for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [89, 90, 91, 92], which enable us to give a comprehensive exploration on the dark matter effects and non-linear electrodynamics effects (or QED effects) in black hole physics.
@@
In this work, we carry out an investigation on the gravitational lensing of Eular-Heisenberg black hole surrounded by PFDM. The gravitational lensing observables provide a avenue to reveal the features of gravitational field and intrinsic properties of black holes through high precision observational data [93]. And an exploration of gravitational lensing observables for Eular-Heisenberg black hole surround by PFDM could enrich our knowledge on the gravitational field for charged black hole systems, especially the influences coming from the dark matter effects and nonlinear electrodynamics effects (or QED effects). Concretely, in the present work, we calculate the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits for Euler-Heisenberg black hole surrounded by PFDM. There are significantly important observables in the gravitational lensing observations, which have been frequently used to test various aspects of black hole physics and gravitational theories in both strong gravitational field limit and weak gravitational field limit. Furthermore, stimulated by the rapid development of black hole shadow since the capture of black hole images by Event Horizon Telescope (EHT) [94, 95, 96], the black hole critical shadow radius for Euler-Heisenberg black hole in the presence of PFDM is also discussed.
@@
This rest of the present work is organized in the following way. Section 2 briefly introduces the spacetime metric generated by Euler-Heisenberg black hole surrounded by PFDM. In section 3, the theoretical approach we employed in calculating the gravitational lensing observables is described in detailed. In section 4, results and discussions on the gravitational lensing observables (gravitational deflection angle of light, time delay of light, trajectory and precession angle of massive object’s bound orbit, black hole shadow critical radius) for Euler-Heisenberg black holes surrounded by PFDM are presented. Summary and perspectives are given in section 5. Throughout this work, the natural unit $G=c=1$ is adopted.
@@ 2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
The Euler-Heisenberg black hole becomes attractive in recent year for its ability to include the nonlinear electrodynamics effects (or QED effects). This kind of black hole can be solved from the gravitational field equation and the Euler-Heisenberg effective action [53], and it is one of the most important realization on the coupling between nonlinear electrodynamics and charged black hole systems.
@@
with the energy-momentum tensor contributed from nonlinear electromagnetic field and dark matter field. To derive a spherically symmetric analytical solution for black hole, the energy-momentum tensor for PFDM and nonlinear electromagnetic field are restricted to the following explicit form in the spherical coordinates [89]
@@
where $M$ is the black hole mass, $Q$ is the black hole electric charge, $a$ is a parameter which measures the strength of non-linear electrodynamics effects (or QED effects). The parameter $\lambda_{\text{DM}}$ denotes the influences of PFDM, which is proportional to the dark matter density. Eventually, by solving the gravitational field equation in the spherically symmetric coordinates, the spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM is
@@
In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric turn to be the Euler-Heisenberg black hole with non-linear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects / QED effect vanishes and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of Euler-Heisenberg black hole can be greatly influenced. It is commonly known that the classical charge black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. However, the Euler-Heisenberg black hole with a nonzero nonlinear electrodynamics parameter $a\neq 0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge. Therefore, in the rest of this work, it is safely to choose black hole parameters to study the gravitational lensing observables, without worrying about the annoying naked singularities. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in the Appendix A.
@@ 3 Theoretical Methods to Obtain Gravitational Lensing Observables
This section gives descriptions on the theoretical methods used to calculate the gravitational lensing observables in our work. The treatments of particle orbits (which are geodesics) are presented in subsection 3.1. Schemes to obtained gravitational deflection angle of light and time delay of light for finite distance light source and observer are introduced in subsection 3.2 - 3.3. Finally, the formulations on the trajectories and precession angle for massive object’s bound orbits around black holes are given in subsection 3.4.
@@ 3.1 Geodesics
In the gravitational filed, massive and massless particles are moving along geodesics in 4 dimensional spacetime. Here we provide a derivation of the reduced geodesic equation, which turned out to be extremely useful in the exploration of gravitational lensing observables. For a spherically symmetric spacetime with the metric form
@@
Here, $V_{\text{eff}}(r)=\frac{f(r)}{2}\big{[}\frac{J^{2}}{r^{2}}+\epsilon\big{]}$ is the effective potential of test particles moving in spherically symmetric gravitational field, and $b\equiv|J/E|$ is the impact parameter in a particle orbit (which is also conserved along this orbit). For massless particles traveling along null geodesics, the parameter $\epsilon$ takes the value $\epsilon=0$. For massive particles traveling along timelike geodesics, the parameter $\epsilon$ takes the value $\epsilon=1$.
@@ 3.2 Gravitational Deflection Angle of Light
The gravitational deflection angle reflect the bending or distortion effects on particle orbit in a gravitational field. In the gravitational lensing observation, when the light source and observer are located at a finite distance region to the central black hole (with the radial coordinate $r=r_{\text{S}}$ and $r=r_{\text{O}}$, respectively), the gravitational deflection angle of light is defined through [97]
@@
The notations $\Psi_{\text{O}}$, $\Psi_{\text{S}}$ denote the angles between light ray propagation direction and the radial direction (which are measured at the observer’s position and source position), and $\Delta\phi_{\text{OS}}$ is the variation of azimuthal angle $\phi$ in the light trajectory. For a spherically symmetric spacetime described by metric expression (7), the angle between light ray propagation direction and the radial direction is uniquely determined by spacetime metric [97], via
@@
In the thin lens approximation, this definition of gravitational deflection angle for finite distance source and observer is illustrated in figure 1. The thin lens approximation assumes that the spacetime is flat everywhere except for the location of thin lens (the central black hole). For the quadrangle depicted in figure 1, the exterior angles at each point are $\Psi_{\text{S}}$, $\alpha$, $\pi-\Psi_{\text{O}}$, $\pi-\Delta\phi_{\text{OS}}$, respectively. Under the thin lens approximation, the sum of exterior angles for the quadrangle equals $2\pi$, which justify the definition of gravitational deflection angle $\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$. However, it should be noted that this definition given by Ishihara et al. is generally valid in gravitational lensing treatments which goes beyond the thin lens approximation, see references [97, 98, 99, 100, 101] for more discussions on this issue.
@@
where we have used the affine parameter $\epsilon=0$ for massless photon, the conserved energy parameter $E=f(r)\frac{dt}{d\lambda}$, conserved angular momentum $J=r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$, and the definition of impact parameter $b\equiv|J/E|$. The plus and minus sign $\pm$ can be determined in the following way. When the photon beams travel along a scattering orbit from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as the azimuthal angle $\phi$ increases, until photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase with the increasing of azimuthal angle.
@@
$\frac{dr}{d\lambda}\bigg{|}_{r=r_{0}}=0\ \ \ \Rightarrow\ \ \ \frac{1}{2}f(r_{0})\frac{J^{2}}{r_{0}^{2}}=\frac{1}{2}E^{2}\ \ \ \Rightarrow\ \ \ b^{2}=\frac{J^{2}}{E^{2}}=\frac{r_{0}^{2}}{f(r_{0})}$ (14)

### v2

@@ Abstract
In this work, we study the gravitating lensing Euler-Heisenberg black hole surrounded by perfect fluid dark matter. This kind of black hole solution enables us to investigate the nontrivial interplay between the dark matter effects and nonlinear electrodynamics effects (or quantum electrodynamics effects) on charged black hole systems. The important observables in gravitational lensings are calculated and discussed in our work, including the gravitational deflection angle of light, time delay of light, precession angle of massive orbit’s bound orbits, and black hole shadow radius. The results indicates that the Euler-Heisenberg black hole with a larger perfect fluid dark matter parameter could greatly reduce the gravitational deflection angle of light, time delay of light and precession angle of massive object’s bound orbit, while the nonlinear electrodynamics effects do not have large influences on these gravitational lensing observables.
Keywords: Euler Heisenberg Black Hole; Gravitational Lensing;
Perfect Fluid Dark Matter; Nonlinear Electrodynamics Effects
@@ 1 Introduction
The equation of state for dark matter in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted large interests [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trial effects and characters for dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and effective spacetime metrics for several kinds of black holes (such as Schwarzschild, RN, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties of black hole system in PFDM medium can be extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
Inspired by the simplicity and success of PFDM model in black hole physics and astronomy, we choose a novel black hole solution in PFDM medium — the Euler-Heisenberg black hole surrounded by PFDM. The Euler-Heisenberg black hole, which is obtained from the Euler-Heisenberg effective action [53, 54, 55, 56], stimulated numerous studies in past few years, for its nontrivial realization on the coupling between nonlinear electrodynamics and gravity systems, especially with the quantum loop-corrections in quantum electrodynamics (QED) [57]. With these desired properties, this kind of black hole is often used to explore the nonlinear electrodynamics in black hole physics, evoking a number of studies on thermodynamic transitions, thermodynamic topology, black hole shadow, gravitational lensing, quasi-normal modes for black hole systems with nonlinear electrodynamics in recent years [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 84, 80, 81, 82, 83, 89, 86, 87, 91, 88, 90, 85]. The nonlinear electrodynamics effects (or QED effects) on gravitation are extremely important issues in high energy scale, and they may have potential influences on the evolution of our universe (especially at the beginning stage), such as the inflationary theory in Big Bang, the dynamic evolution of dark matter and dark energy, and the cosmological accelerating expansion of universe. Therefore, it is interesting to study the interplay between dark matter and nonlinear electrodynamics in gravitational systems. The Euler-Heisenberg black hole surrounded by PFDM could provide us a simply way to investigate their influences on charged black holes. Recently, several analytical solution for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [92, 93, 94, 95], which enable us to give a comprehensive exploration on the interplay dark matter effects and nonlinear electrodynamics effects in black hole systems.
@@
In this work, we carry out an investigation on the gravitational lensing of Eular-Heisenberg black hole surrounded by PFDM. The gravitational lensing observables provide an avenue to reveal the features of gravitational field and intrinsic properties of black holes through high precision observational data [96, 97]. An exploration of gravitational lensing observables for Eular-Heisenberg black hole surround by PFDM could enrich our knowledge on the gravitational field for charged black hole systems, especially with the influences coming from the dark matter effects and nonlinear electrodynamics effects (or QED effects). Concretely, in the present work, we calculate the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits for Euler-Heisenberg black hole surrounded by PFDM. There are significantly important observables in the gravitational lensing observations, which have been frequently used to test various aspects of black hole physics and gravitational theories in both strong gravitational field limit and weak gravitational field limit. Furthermore, stimulated by the rapid development of black hole shadow researches since the capture of black hole images by Event Horizon Telescope (EHT) [98, 99, 100], the black hole critical shadow radius for Euler-Heisenberg black hole in the presence of PFDM is also discussed in our work.
@@
This rest of the present work is organized in the following way. Section 2 briefly introduces the spacetime metric generated by Euler-Heisenberg black hole surrounded by PFDM. In section 3, the theoretical approach employed in calculating the gravitational lensing observables is described in details. In section 4, numerical results and discussions on the gravitational lensing observables (gravitational deflection angle of light, time delay of light, trajectory and precession angle of massive object’s bound orbit, black hole shadow critical radius) for Euler-Heisenberg black holes surrounded by PFDM are presented. Summary and perspectives are given in section 5. Throughout this work, the natural unit $G=c=1$ is adopted.
@@ 2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
The Euler-Heisenberg black hole becomes attractive in recent year for its ability to include the nonlinear electrodynamics effects (or QED effects). This kind of black hole solution can be solved from the gravitational field equation and the Euler-Heisenberg effective action [53], and it is one of the most important realization on the coupling between nonlinear electrodynamics and charged black hole systems.
@@
with the energy-momentum tensor contributed from nonlinear electromagnetic field and dark matter field. To derive a spherically symmetric analytical solution for black hole, the energy-momentum tensors for PFDM and nonlinear electromagnetic field are restricted to the following explicit form in the spherical coordinate [92]
@@
where $M$ is the black hole mass, $Q$ is the black hole electric charge, $a$ is a parameter which measures the strength of nonlinear electrodynamics effects (or QED effects). The parameter $\lambda_{\text{DM}}$ denotes the influences of PFDM, which is proportional to the dark matter density. Eventually, by solving the gravitational field equation in the spherical coordinates, the spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM is
@@
In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric represents an Euler-Heisenberg black hole with nonlinear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects vanish and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of charged black hole can be greatly influenced. It is commonly known that the classical charged black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. However, the Euler-Heisenberg black hole with a nonzero nonlinear electrodynamics parameter $a>0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge (see Appendix A). Therefore, in the rest of this work, it is safely to change black hole parameters in the calculation of gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, without worrying about the annoying naked singularities. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in the Appendix A.
As an theoretical investigation, we only consider the dark matter field and nonlinear electromagnetic field in this work. The influences from other complex matter fields (such as disk structure in galaxies composed of luminous stars, dusts and plasma in galaxies) are not included in our work. Since the energy-momentum tensor of the dark matter fluid is spherically symmetric, they can form a spherically symmetric halo structure around the central supermassive black hole. In this work, we assume that the event horizon of Euler-Heisenberg black hole (surrounded by PFDM) $r_{H}$, the scale of dark matter halo $r_{\text{halo}}$, and the impact parameter for particle orbits $b$ satisfy
$r_{H}\ll b\ll r_{\text{halo}}$ (7)
Furthermore, it should be mentioned that the PFDM model which gives rise to the black hole solution in (6) is a very simplified perfect fluid model. In such simplified model, the energy-momentum tensor of PFDM contains only one free parameter $\lambda_{\text{DM}}$, so the energy density $\rho$ and pressure $P$ of the isotropic dark matter fluid are directly linked with $\lambda_{\text{DM}}$ via expression (4). The spacetime metric of black holes in this simple dark matter model satisfies $g_{tt}\times g_{rr}=-1$. Recently, some studies have explored more complex perfect fluid models for dark matter. In these dark matter fluid models, the energy-momentum tensor of dark matter fluid contains the energy density $\rho$ and the pressures in radial direction and tangential direction (parameterized by $P_{r}$ and $P_{t}$), and they are connected through the equation of state for dark matter fluids [101, 102, 103, 104]. Especially, some researchers are interested in the Tolman-Oppenheimer-Volkoff (TOV) equation of state for dark matter fluids [104, 105]. The charged black hole systems solved form these dark matter models are more complex, and the $g_{tt}\times g_{rr}=-1$ relation for spacetime metric is no longer valid. The studies of more complex charged black hole solutions from dark matter model is beyond the scope of our present work.
@@ 3 Theoretical Methods to Obtain Gravitational Lensing Observables
This section gives descriptions on the theoretical methods used to calculate the gravitational lensing observables in our work. The treatments of particle orbits (which are geodesics) are presented in subsection 3.1. Schemes to obtained gravitational deflection angle of light and time delay of light for finite distance light source and observer are introduced in subsections 3.2 - 3.3. Finally, the formulations on the trajectories and precession angle for massive object’s bound orbits around black holes are given in subsection 3.4.
@@ 3.1 Geodesics
In the gravitational filed, massive and massless particles are moving along geodesics in 4 dimensional spacetime. Here we provide a derivation of the reduced geodesic equation, which turns out to be extremely useful in the exploration of gravitational lensing observables. For a spherically symmetric spacetime with the metric form
@@
Here, $V_{\text{eff}}(r)=\frac{f(r)}{2}\big{[}\frac{J^{2}}{r^{2}}+\epsilon\big{]}$ is the effective potential of test particles moving in a spherically symmetric gravitational field, and $b\equiv|J/E|$ is the impact parameter in a particle orbit (which is also conserved along this orbit). For massless particles traveling along null geodesics, the parameter $\epsilon$ takes the value $\epsilon=0$. For massive particles traveling along timelike geodesics, the parameter $\epsilon$ takes the value $\epsilon=1$.
@@ 3.2 Gravitational Deflection Angle of Light
The gravitational deflection angle reflects the bending or distortion effects on particle orbit in a gravitational field. In the gravitational lensing observation, when the light source and observer are located at a finite distance region to the central black hole (with the radial coordinate $r=r_{\text{S}}$ and $r=r_{\text{O}}$, respectively), the gravitational deflection angle of light is defined through [106]
@@
The notations $\Psi_{\text{O}}$, $\Psi_{\text{S}}$ denote the angles between light ray propagation direction and the radial direction (which are measured at the observer’s position and source position), and $\Delta\phi_{\text{OS}}$ is the variation of azimuthal angle $\phi$ in the light trajectory. For a spherically symmetric spacetime described by metric (8), the angle between light ray propagation direction and the radial direction is uniquely determined by spacetime metric [106], via
@@
The definition of gravitational deflection angle of light in expression (11) for finite distance light source and observer can be understood from an instructive way using thin lens approximation. In the thin lens approximation, the gravitational deflection angle for finite distance light source and observer is illustrated in figure 1. The thin lens approximation assumes that the spacetime is flat everywhere except for the location of gravitational thin lens (the central black hole position). For the quadrangle depicted in figure 1, the exterior angles at each point are $\Psi_{\text{S}}$, $\alpha$, $\pi-\Psi_{\text{O}}$, $\pi-\Delta\phi_{\text{OS}}$, respectively. Under the thin lens approximation, the sum of exterior angles for the quadrangle equals $2\pi$, which justify the definition of gravitational deflection angle $\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$. However, it should be noted that this definition in expression (11) given by Ishihara et al. is not restricted in the thin lens approximation. It is a consistent and proper definition of gravitational deflection angle that are hold in more generally gravitational lensing treatments (which can go beyond the thin lens approximation), see references [106, 107, 108, 109, 110] for more discussions on this issue.
@@
where we have used the affine parameter $\epsilon=0$ for massless photon, the conserved energy parameter $E=f(r)\frac{dt}{d\lambda}$, conserved angular momentum $J=r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$, and the definition of impact parameter $b\equiv|J/E|$. The plus and minus sign $\pm$ can be determined in the following way. When the photon beams travel along a scattering orbit from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as the azimuthal angle $\phi$ increases, until photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase with the increasing of azimuthal angle. This we derive the following relation
@@
$\frac{dr}{d\lambda}\bigg{|}_{r=r_{0}}=0$$\Rightarrow$ $\frac{1}{2}f(r_{0})\frac{J^{2}}{r_{0}^{2}}=\frac{1}{2}E^{2}$ (15)
$\Rightarrow$ $b^{2}=\frac{J^{2}}{E^{2}}=\frac{r_{0}^{2}}{f(r_{0})}$